OpenStudy (anonymous):

solve for y: x - 1/9y = -1

OpenStudy (anonymous):

you mean -y/9 or is y supposed to be on the bottom?

OpenStudy (anonymous):

I'm guessing y is on top so I'll solve it that way.

OpenStudy (anonymous):

x - 1*y/9 = -1 -1*y/9 = -1 -x y/9 = 1 + x y = 9*(1+x) y = 9 + x

Directrix (directrix):

No diagram?

OpenStudy (anonymous):

y = 9 + 9x

OpenStudy (anonymous):

sorry it's late!

OpenStudy (anonymous):

direct, I'm sure they want it step by step if they're learning

OpenStudy (anonymous):

I thought i attached it

OpenStudy (anonymous):

Yep, that means y is on top.

OpenStudy (anonymous):

thanks

Directrix (directrix):

If the equation is x - (1/9)y = -1 then I cranked out y = 9 + 9x.

Directrix (directrix):

That would be A but I may have messed up. Will check.